### v1

@@ Abstract
New boron abundances or upper limits have been determined for 8 early-B stars in the young Galactic open cluster NGC 3293, using ultraviolet spectra obtained by the Hubble Space Telescope Cosmic Origins Spectrograph. With previous observations, there are now 18 early-B stars in this cluster with boron measurements. Six of the newly observed stars have projected rotational velocities greater than 200 km/s, allowing new constraints on rotationally driven mixing in main-sequence stars. When comparing to synthetic model populations, we find that the majority of our sample stars agree well with the predicted trends of stronger boron depletion for larger rotation and for larger mass or luminosity. Based on those, a smaller than the canonical rotational mixing efficiency,( vs the more standard value of 0.033), appears to be required. However, our five most slowly rotating stars are not well explained by rotational mixing, and we speculate that they originate from binary mergers.
@@ 1 Introduction
The dominant energy source in core hydrogen burning massive stars is the fusion of protons to He. Since massive stars use the CNO-cycle to perform this, most of the available carbon and oxygen in the core is converted into nitrogen during this process. At the same time, the pp-chains are also operating. While they contribute little to the energy production in massive stars, the corresponding proton capture reactions leads to the destruction of the light elements, lithium, beryllium and boron, in most of the stellar interior.
@@
However, refined observations cast doubts on this explanation. Hunter et al. (2008b), identified a population of slowly rotating early B dwarfs in the LMC which showed a significant surface nitrogen enrichment. A similar population was also found in the LMC O star population (Rivero González et al., 2012; Grin et al., 2017). At the same time, it became clear that a typical population of massive main sequence single stars must be expected to contain a significant fraction (%) of binary products, i.e., mass gainers in mass transfer systems or mergers, (Sana et al., 2012; de Mink et al., 2014).
In this paper, we pursue the idea that boron abundance measurements may have a strong potential to discriminate between the suggested enrichment channels. Proffitt et al. (2016) discussed observations of the B iii resonance line in 10 early-B stars in the young open cluster NGC 3293, and concluded that modest boron depletion was seen, perhaps consistent with the predictions of the models of Brott et al. (2011a), in which rotational mixing was calibrated using observations of nitrogen abundances in massive main sequence stars of the LMC. We requested observations of ten additional stars, and obtained useful data for 8 of these, as part of HST program 14673, with a focus on more rapidly rotating stars lower on the main-sequence. These data provide us with a unique sample of 18 early B dwarfs with information on their boron and nitrogen abundances in an open star cluster, i.e., with roughly known and equal ages.
@@ 2 UV observations
Targets for both programs were selected from Dufton et al. (2006), based on their location in the HR diagram as well as their rotational velocity. For the new program, we gave preference to more rapidly rotating stars. After the UV observations were obtained, we found that ESL 016 was a much narrower lined star than we had expected. While Dufton et al. (2006) had adopted the correct photometric information for this star, the FLAMES-Giraffe Spectrograph observation that they associated with it appears to have actually been of the fainter, (), star NGC 3293 B 198, (apparently Gaia EDR3 5351447238427079936), (Balona, 1994), which is located about 2” away from ESL 016. Fortunately, the ESO archive does contain three similar FEROS observations of the brighter star, under the name V 404 Car, taken in 2016 by F. Rodler as part of ESO program 079.A-9008, and we were able to use these spectra for our analysis.
@@
For one of the targets, ESL 020, a bad guide star acquisition caused a mis-pointed observation with the target on the edge of the aperture and only a very faint spectrum visible. Since the program was over 90% complete, this visit was not eligible for a repeat. In addition, for ESL 025, the observed UV spectrum appears to be a composite that does not correspond to the optical spectrum and could not be usefully fit. While this star does have a companion about 1.6 magnitudes fainter at a distance of about 2.3”, that star should have been well out of the 2.5” diameter COS aperture, and the 2D spectral image shows no evidence for multiple components. As a result, these two stars will be left out of our analysis.
@@
3 Determination of Stellar Parameters from Optical Spectra
@@
The earlier analysis of this cluster by Dufton et al. (2006) and Hunter et al. (2009) primarily relied upon H and He line profiles (H, H, and He i 4026Å) to determine and , while for some stars where the He lines were not useful, parameters were estimated based on the spectral type. For the most narrow lined stars, the Si ionization ratios were also used to constrain the effective temperature.
@@
To determine revised stellar parameters, we reanalyzed the same optical spectra used by Dufton et al. (2006). These observations were originally described in Evans et al. (2005). For the brighter stars, (ESL 002 through ESL 019), in our sample, these spectra were obtained with the Fibre-Fed Extended Range Optical Spectrograph (FEROS) at the Max Planck Gesellschaft MPG/ESO telescope, with ), while for the remainder of the stars, the Fibre Large Array Multi-Element Spectrograph (FLAMES) instrument at the Very Large Telescope (VLT) was used with the FLAMES-Giraffe Spectrograph ().
For four of the narrower lined stars in our sample (ESL 002, 003, 007, and 010), we applied the iterative scheme and updated TLUSTY models presented in Bragança et al. (2019) to redetermine , , abundances of silicon and oxygen, and other stellar parameters. For the other cluster stars in our sample, we then froze the Si and O abundances to the average value determined for these benchmark stars, and then redetermined the stellar parameters based on four spectral regions, the CNO lines near 4650Å, H, He i 4471 Å, and the Si iii triplet at 4552 - 4575 Å.
@@
The procedure used was as follows: The procedure used was as follows: 1) We started with the initial and from Dufton et al. (2006). 2) We refit the CNO region near 4650 Å to adjust , while keeping fixed. 3) We adjusted to fit the H line while keeping fixed. 4) We again adjusted to fit the 4471 Å He i line while keeping fixed. 5) Using the from step 3 and the average of the values from steps 2 and 4, we fitted the Si iii triplet while keeping the Si abundance fixed at the mean abundance found for the four benchmark stars 6) Finally, for sample stars with 24 000 K, we checked the fit to He ii 4686 Å to verify the final parameters.
The resultant parameters are listed in Table 2 and for comparison we also list the parameters derived by Dufton et al. (2006). We find that the present effective temperature and surface gravity estimates are higher than in Dufton et al. (2006), especially for the hotter stars classified as B1 or B1.5. This is consistent with Proffitt et al. (2016) who argued that these higher values, derived from the Si ii/Si iii ionization balance, are to be preferred to the lower values, that were derived from the Si iii/Si iv ionization balance.
@@
Four of our sources, ESL 004, 007, 008, and 019, have HST G430L spectrophotometry (Fitzpatrick et al., 2019) covering the important Balmer jump region, that is sensitive to both effective temperature and surface gravity. These data were used to check the present solutions for these stars by finding the model that best reproduces the G430L spectral energy distribution (SED). We used the Levenberg-Marquardt algorithm, (Markwardt, 2009), and a grid of TLUSTY models for each star, normalized to the observed spectrum in the -band, with extinction applied according to the observed and model measurements and appropriate rotational and radial velocity corrections, and then re-binned to the G430L resolution and pixel size. The results are illustrated in Figure 1 and summarized in Table 3.
@@
Inspection of Figure 1 confirms that the parameters derived from our spectroscopic analysis of the high resolution optical data also provide an excellent match to observed SEDs, except for star ESL 007. However close inspection of the Balmer jump of that star reveals the presence of a weak secondary emission jump, suggesting this is a Be star (Cochetti et al., 2020), an hypothesis supported by the presence of previously overlooked weak emission in the core of the H profile. While weak narrow nebular H emission is present for most of the sources discussed here, that of ESL 007 is clearly broader than nebular emission. The spectroscopic parameters are therefore to be preferred over the SED results for ESL 007. Comparing results for the other three stars it is clear that the higher effective temperatures derived for the B1/B1.5 stars using the present spectroscopic and SED approaches are indeed appropriate for this spectral type.

### v2

@@ Abstract
New boron abundances or upper limits have been determined for 8 early-B stars in the young Galactic open cluster NGC 3293, using ultraviolet spectra obtained by the Hubble Space Telescope Cosmic Origins Spectrograph. With previous observations, there are now 18 early-B stars in this cluster with boron measurements. Six of the newly observed stars have projected rotational velocities greater than 200 km/s, allowing new constraints on rotationally driven mixing in main-sequence stars. When comparing to synthetic model populations, we find that the majority of our sample stars agree well with the predicted trends of stronger boron depletion for larger rotation and for larger mass or luminosity. Based on those, a smaller than the canonical rotational mixing efficiency, ( vs the more standard value of 0.033), appears to be favored. In addition, the five mostly slowly rotating stars, when considered as a group, tend to show more boron depletion than expected from rotational mixing, and we speculate that most or all of these originate from binary mergers.
@@ 1 Introduction
The dominant energy source in core hydrogen burning massive stars is the fusion of protons to 4He. Since massive stars use the CNO-cycle to perform this, most of the available carbon and oxygen in the core is converted into nitrogen during this process. At the same time, the pp-chains are also operating. While they contribute little to the energy production in massive stars, the corresponding proton capture reactions leads to the destruction of the light elements, lithium, beryllium and boron, in most of the stellar interior.
@@
However, refined observations cast doubts on this explanation. Hunter et al. (2008b), identified a population of slowly rotating early B dwarfs in the LMC which showed a significant surface nitrogen enrichment. A similar population was also found in the LMC O star population (Rivero González et al., 2012; Grin et al., 2017). Similar discrepancies have been found in Galactic stars. Some well studied stars that are known to be intrinsically slow rotators nevertheless show evidence of mixing, (e.g., Sco as discussed by Keszthelyi et al., 2021), while Cazorla et al. (2017) finds fast rotators without the expected nitrogen enrichment, and Markova et al. (2014, 2018), discuss slowly rotating nitrogen rich stars. At the same time, it became clear that a typical population of massive main sequence single stars must be expected to contain a significant fraction (%) of binary products, i.e., mass gainers in mass transfer systems or mergers, (Sana et al., 2012; de Mink et al., 2014).
In this paper, we pursue the idea that boron abundance measurements may have a strong potential to discriminate between the suggested enrichment channels. Proffitt et al. (2016) discussed observations of the B III resonance line in 10 early-B stars in the young open cluster NGC 3293, and concluded that modest boron depletion was seen, and that for most stars the depletion was slightly less efficient than predicted by the models of Brott et al. (2011a), in which rotational mixing was calibrated using observations of nitrogen abundances in massive main sequence stars of the LMC. In particular they noted that three of the ten stars appeared to be slightly boron rich relative to the expectations of these models. We requested observations of ten additional stars, and obtained useful data for 8 of these, as part of HST program 14673, with a focus on more rapidly rotating stars lower on the main-sequence. These data provide us with a unique sample of 18 early B dwarfs with information on their boron and nitrogen abundances in an open star cluster, i.e., with roughly known and equal ages.
@@ 2 UV observations
Targets for both programs were selected from Dufton et al. (2006), based on their location in the HR diagram as well as their rotational velocity. For the new program, we gave preference to more rapidly rotating stars. An additional important criterion was cluster membership, as implied by both photometric and spectroscopic surveys of the cluster (Baume et al., 2003; Dufton et al., 2006). Subsequently, Göppl & Preibisch (2022) have analysed data for this cluster and confirm cluster membership for all sources considered here, barring possibly ESL 028, that they comment on as a possible 2 background star based on its parallax measurement. This star has a RUWE value of 1.50, which is is out-of-family with other cluster members that have values close to 0.9, and based on its overall properties it is still designated a cluster member by Morel et al. (2022).
One might consider using radial velocity information to further refine cluster membership probability. However as noted by Feast (1958) there are a significant number of radial velocity variables in the cluster. Those data suggest that much of this variability is associated with Cep variables, which are common in this cluster, (see Table 1 for known variables), given the temperature of its turn-off stars. Moreover a preliminary analysis of ESO archival data for selected Cep members confirms significant apparent velocity shifts occur on short timescales, comparable to their periods. However we agree with Feast (1958) that ESL 008 is a potential SB1 system with a peak-to-peak velocity of 90 km/s, cluster membership implying a minimum semi-amplitude velocity of 70 km/s. Note that while Morel et al. (2022) suggest this source to be an SB2 system, their analysis is based on a single VLT/UVES spectrum. However comparing their data with the 2.2m/FEROS data of Dufton et al. (2006) we find no evidence of a secondary in either spectrogram.
After the UV observations were obtained, we found that ESL 016 was a much narrower lined star than we had expected. While Dufton et al. (2006) had adopted the correct photometric information for this star, the FLAMES-Giraffe Spectrograph observation that they associated with it appears to have actually been of the fainter, (), star NGC 3293 B 198, (apparently Gaia EDR3 5351447238427079936), (Balona, 1994), which is located about 2” away from ESL 16. Morel et al. (2022), appears to have adopted the same mis-identification. Fortunately, the ESO archive does contain three similar FEROS observations of the brighter star, under the name V 404 Car, taken in 2016 by F. Rodler as part of ESO program 079.A-9008, and we were able to use these spectra for our analysis.
@@
For one of the targets, ESL 20, a bad guide star acquisition caused a mis-pointed observation with the target on the edge of the aperture and only a very faint spectrum visible. Since the program was over 90% complete, this visit was not eligible for a repeat. In addition, for ESL 25, the observed UV spectrum appears to be a composite that does not correspond to the optical spectrum and could not be usefully fit. While this star does have a companion about 1.6 magnitudes fainter at a distance of about 2.3”, that star should have been well out of the 2.5” diameter COS aperture, and the 2D spectral image shows no evidence for multiple components. As a result, these two stars will be left out of our analysis.
3 Determination of Stellar Parameters
3.1 Discussion of Literature Results
@@
An extensive spectroscopic analysis of B stars in NGC 3293 was included as part of the FLAMES survey (Evans et al., 2005), and the resulting stellar parameter and abundance determinations were presented in Dufton et al. (2006) and Hunter et al. (2009). This work primarily relied upon H and He line profiles (H, H, and He I 4026Å) to determine and , while for some stars where the He lines were not useful, parameters were estimated based on the spectral type. For the most narrow lined stars, the Si ionization ratios were also used to constrain the effective temperature.
A recent paper by Morel et al. (2022) reanalyzed a subset of these stars and determined stellar parameters from a global fit to to several spectral regions that included H, H, as well as various lines of He I, Si II, Si III, O II that put useful constraints on the stellar parameters. They then determined abundances from selected lines of individual elements, including C II 4267 and N II 4630. Morel et al. (2022), however, did not attempt to measure the value of the microturbulence, , instead using fixed values of 2 or 5 km/s depending on stellar parameters.
For the stars we consider, we show the results for stellar parameters, carbon, and nitrogen abundances found by both Hunter et al. (2009) and Morel et al. (2022) in Table 2, and will discuss their findings in the context of our new results below. For those results presented by Hunter et al. (2009) where the parameters were estimated from the spectral type, we mark the with a colon to indicate the greater uncertainty in the parameters.
| ESL# | Dufton et al. (2006) and Hunter et al. (2009)aaFor most stars Hunter et al. (2009) adopted the stellar parameters from Dufton et al. (2006), but slightly updated the parameters for a few stars in addition to deriving abundances. values which Dufton et al. (2006) derived completely or partially from estimated spectral types have larger uncertainties and are marked with a colon. | Morel et al. (2022) | ||||||||||
| C | N | C | N | |||||||||
| 002 | 23200 | 2.80 | 15 | 100 | ||||||||
| 003 | 20500 | 2.75 | 13 | 80 | ||||||||
| 004 | 22700 | 3.13 | 13 | 105 | ||||||||
| 005 | 21500: | 3.05 | 14 | 195 | ||||||||
| 006 | 21500: | 3.15 | 15 | 200 | ||||||||
| 007 | 22600 | 3.10 | 11 | 65 | ||||||||
| 008 | 21500: | 3.25 | 12 | 140 | ||||||||
| 010 | 21450 | 3.20 | 11 | 70 | 24493 | 3.61 | 5 | 41 | ||||
| 012 | 21500: | 3.30 | 11 | 100 | 24444 | 3.70 | 5 | 83 | ||||
| 015 | 25000: | 3.80 | 260 | |||||||||
| 019 | 25000: | 3.85 | 6 | 120 | 24380 | 3.63 | 5 | 107 | ||||
| 023 | 20500: | 3.40 | 6 | 160 | 21730 | 3.71 | 2 | 154 | ||||
| 024 | 20500: | 3.50 | 6 | 135 | 21700 | 3.69 | 2 | 132 | ||||
| 025 | 21100 | 3.70 | 5 | 215 | 20460 | 3.69 | 2 | 192 | ||||
| 028 | 19400 | 3.65 | 7 | 215 | 20429 | 3.86 | 2 | 165 | ||||
| 030 | 19800 | 3.70 | 4 | 205 | 19675 | 3.80 | 2 | 216 | ||||
| 031 | 17400 | 3.45 | 12 | 230 | 20322 | 3.86 | 2 | 257 | ||||
| 038 | 20600 | 3.95 | 1 | 235 | 20216 | 3.94 | 2 | 229 | ||||
3.2 New Analysis of Optical Spectra
@@
To determine revised stellar parameters, we reanalyzed the same optical spectra used by Dufton et al. (2006). These observations were originally described in Evans et al. (2005). For the brighter stars, (ESL 02 through ESL 19), in our sample, these spectra were obtained with the Fibre-Fed Extended Range Optical Spectrograph (FEROS) at the Max Planck Gesellschaft MPG/ESO telescope, with and coverage between 3600 and 9200 Å), while for the remainder of the stars, the Fibre Large Array Multi-Element Spectrograph (FLAMES) instrument at the Very Large Telescope (VLT) was used with the FLAMES-Giraffe Spectrograph (). Giraffe settings HR02, (cenwave 3958), HR03 (4124), HR04 (4297), HR05 (4471), HR06 (4656), and HR14 (6515) were used providing coverage from 3850–4475Å and 6380–6620Å were used.
In this study, we selected the four sample stars in NGC 3293, having the lowest , as benchmark stars for which we could analyze a large number of spectral lines. The selected stars were: ESL 02, 03, 07, and 10. We then used the wrapper Stellar Spectral Synthesis Suite (S4, Bragança et al. (2019)) to determine their , , the abundances of silicon and oxygen, as well as the projected rotational velocity , the radial-tangential macroturbulent velocity , and the microturbulent velocity . The iterative scheme of S4 interpolates in a grid of non-LTE model atmospheres computed with TLUSTY and using updated model atoms of oxygen, carbon, silicon and neon. The iteration begins with initial guesses of the stellar parameters: , , , , , and O and Si abundances (initially adopted as the solar values obtained by Asplund et al. 2009). The ionization balance of silicon (based on Si II lines 4128, 4130, 6347 and 6371 Å and Si III) lines 4552, 4567, 4574 and 4716 Å) is used to constrain the , while the wings of the Hydrogen lines H and H are used to define . The selected absorption profiles are fitted independently and the best fits between synthesis and observations are obtained by minimizations. The silicon abundances are derived for a range of microturbulent velocity values, and the selected is the one that corresponds to the lowest dispersion in the abundance distribution. is allowed to vary in order to obtain the best fit for every profile and is included as an additional broadening parameter needed to fit the wings of metal lines. Once the stellar parameters are defined, one more iteration produces the final abundances of oxygen and silicon.
For the rest of our sample, which present shallower and blended lines due to higher values of , the iterative method described above, which was used for the sharp lined stars, could not be directly applied. We then determined the parameters of these stars based on best fits of four spectral regions: 4630-4665 Å, containing CNO lines; 4280-4410 Å, containing the line; 4450-4500 Å, containing He I 4471 Å line, and 4540-4585 Å, containing the Si III triplet at 4552 - 4575 Å. For obtaining the best fits between models and observations we adopted the average Si and O abundances derived for the benchmark stars.
@@
The procedure used was as follows: 1) We started with the initial and from Dufton et al. (2006) 2) We fit the CNO region near 4650 Å to adjust , while keeping fixed. 3) We adjusted to fit the H line while keeping fixed. 4) We again adjusted to fit the 4471 Å He I line while keeping fixed. 5) Using the from step 3 and the average of the values from steps 2 and 4, we fitted the Si III triplet while keeping the Si abundance fixed at the mean abundance found for the four benchmark stars and varying the microturbulent velocity. 6) Finally, for sample stars with 24 000 K, we checked the fit to He II 4686 Å to verify the final parameters. The lines of the rapidly rotating stars are dominated by rotational broadening, making it difficult to constrain the macroturbulent velocity. We fixed to 15 km/s for these stars, except for ESL 04, which required =30 km/s in order to obtain better fits to the profiles. In addition, values were allowed to vary in order to obtain the best fit for the Si III triplet and the blended region at 4650 Å.
The resultant parameters are listed in Table 3 and, for ease of comparison, we also repeat the parameters from Table 2 derived by Morel et al. (2022) where they are available, and from Hunter et al. (2009) or Dufton et al. (2006) otherwise. We find that the present effective temperature and surface gravity estimates are higher than in Dufton et al. (2006), especially for the hotter stars classified as B1 or B1.5. This is consistent with Proffitt et al. (2016) who argued that these higher values, derived from the Si II/Si III ionization balance, are to be preferred to the lower values, that were derived from the Si III/Si IV ionization balance. After excluding ESL 25, for which we could not fit the UV spectrum, for the remaining nine stars in common with Morel et al. (2022), agreement is generally quite good, with a mean difference, (this paper minus Morel), in temperature of K, and mean differences in of less than 0.1 dex. Note that neither Dufton et al. (2006), nor Hunter et al. (2009), nor Morel et al. (2022), included macroturbulent broadening in their analysis. As a result, for several of the more luminous and slower rotating stars where we have included significant macro-turbulent broadening in our fits, we find noticeably smaller projected rotational velocities than was found previously. However, for the majority of the stars, macroturbulence does not make a significant contribution to the line profiles.
| NGC 3293 | This work | Independent Results | ||||||||
| ESL No. | RV | A(Si) | A(O) | |||||||
| 002aaStars fit using the full procedure of Bragança et al. (2019). | 14 | 65 | 42 | 23200bbHunter et al. (2009) | 2.80 | |||||
| 003aaStars fit using the full procedure of Bragança et al. (2019). | 12 | 54 | 27 | 20500bbHunter et al. (2009) | 2.75 | |||||
| 004 | 15 | 95 | 30 | mean | mean | 22700bbHunter et al. (2009) | 3.13 | |||
| 005 | 12 | 180 | 15 | mean | mean | 21500ccDufton et al. (2006) | 3.05 | |||
| 006 | 15 | 200 | 15 | mean | mean | 21500ccDufton et al. (2006) | 3.15 | |||
| 007aaStars fit using the full procedure of Bragança et al. (2019). | 12 | 28 | 26 | 22600bbHunter et al. (2009) | 3.25 | |||||
| 008 | 12 | 130 | 15 | mean | mean | 21500ccDufton et al. (2006) | 3.25 | |||
| 010aaStars fit using the full procedure of Bragança et al. (2019). | 11 | 35 | 30 | 24493ddMorel et al. (2022) | 3.61 | |||||
| 012 | 12 | 90 | 15 | mean | mean | 24444ddMorel et al. (2022) | 3.70 | |||
| 015 | 6 | 260 | 15 | mean | mean | 25000ccDufton et al. (2006) | 3.80 | |||
| 016 | 10 | 55 | 15 | mean | mean | |||||
| 019 | 8 | 120 | 15 | mean | mean | 24380ddMorel et al. (2022) | 3.63 | |||
| 023 | 5 | 170 | 15 | mean | mean | 21730ddMorel et al. (2022) | 3.71 | |||
| 024 | 5 | 160 | 15 | mean | mean | 21700ddMorel et al. (2022) | 3.69 | |||
| 025 | 5 | 215 | 15 | mean | mean | 20460ddMorel et al. (2022) | 3.69 | |||
| 028 | 4 | 200 | 15 | mean | mean | 20429ddMorel et al. (2022) | 3.86 | |||
| 030 | 6 | 220 | 15 | mean | mean | 19675ddMorel et al. (2022) | 3.80 | |||
| 031 | 6 | 230 | 15 | mean | mean | 20322ddMorel et al. (2022) | 3.86 | |||
| 038 | 6 | 235 | 15 | mean | mean | 20216ddMorel et al. (2022) | 3.94 | |||
@@
Four of our sources, ESL 04, 07, 08, and 19, have HST G430L spectrophotometry (Fitzpatrick et al., 2019) covering the important Balmer jump region, that is sensitive to both effective temperature and surface gravity. These data were used to check the present solutions for these stars by finding the model that best reproduces the G430L spectral energy distribution (SED). We used the Levenberg-Marquardt algorithm, (Markwardt, 2009), and a grid of TLUSTY models for each star, normalized to the observed spectrum in the -band, with extinction applied according to the observed and model measurements and appropriate rotational and radial velocity corrections, and then re-binned to the G430L resolution and pixel size. The results are illustrated in Figure 1 and summarized in Table 4.
@@
Inspection of Figure 1 confirms that the parameters derived from our spectroscopic analysis of the high resolution optical data also provide an excellent match to observed SEDs, except for star ESL 007. However close inspection of the Balmer jump of that star reveals the presence of a weak secondary emission jump, suggesting this is a Be star (Cochetti et al., 2020), an hypothesis supported by the presence of previously overlooked weak emission in the core of the H profile. While weak narrow nebular H emission is present for most of the sources discussed here, that of ESL 007 is clearly broader than nebular emission.
The spectroscopic parameters are therefore to be preferred over the SED results for ESL 007. Comparing results for the other three stars it is clear that the higher effective temperatures derived for the B1/B1.5 stars using the present spectroscopic and SED approaches are indeed appropriate for this spectral type.